OpenStudy (anonymous):

Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number?

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

3x - 7 = 5x + 9 Can you solve that? :)

OpenStudy (anonymous):

- 5x from both sides right?

OpenStudy (kamibug):

Right! ^.^ Next?

OpenStudy (anonymous):

3x-7-5x = 5x+9-5x?

OpenStudy (kamibug):

Right. What does that give you?:)

OpenStudy (anonymous):

2x-7=9

OpenStudy (kamibug):

The 2x should be negative :) ... 3x - 5x = -2x

OpenStudy (anonymous):

ooh ok. so now i divide right

OpenStudy (anonymous):

i mean subtract 7 lol

OpenStudy (kamibug):

Add 7 ;D

OpenStudy (anonymous):

k why add?

OpenStudy (anonymous):

i got 5=9 now

OpenStudy (kamibug):

When you add 7 to both sides you should get this... -2x - 7 = 9 -2x - 7 + 7 = 9 + 7 -2x = 16 Now we divide both sides by -2 to find x :) Get it?

OpenStudy (anonymous):

-8

OpenStudy (kamibug):

Correct!

OpenStudy (anonymous):

Thank yooouuu

OpenStudy (kamibug):

Sure ~~(^_^)~~

OpenStudy (anonymous):

can u check one i did?

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

Okay. ^.^

OpenStudy (anonymous):

2(2w-6)=6(w-6)+14 I got w=5

OpenStudy (kamibug):

Yay! It's right. ;D

OpenStudy (anonymous):

Okay ty :)